# UAR-NVC: A Unified AutoRegressive Framework for Memory-Efficient Neural Video Compression

### v1

@@ V CONCLUSION AND FUTURE WORK
Although the framework has been introduced, several aspects require further improvement. Since the complexity of fitting videos varies significantly based on scene dynamics, future work should explore dynamic clip partitioning in the frame grouping process to better adapt to different motion patterns. In addition, our framework currently relies on existing INR models for clip processing, and the RD performance of UAR-NVC with small GOPs is not yet optimal. A more carefully designed and optimized INR model architecture could further enhance performance. Furthermore, our current approach to model reference compression and reference initialization is relatively simplistic. Future work could focus on refining these two modules to further improve efficiency and rate-distortion performance.

### v2

@@ V CONCLUSION AND FUTURE WORK
Although the framework has been introduced, several aspects require further improvement. Since the complexity of fitting videos varies significantly based on scene dynamics, future work should explore dynamic clip partitioning in the frame grouping process to better adapt to different motion patterns. In addition, our framework currently relies on existing INR models for clip processing, and the RD performance of UAR-NVC with small GOPs is not yet optimal. A more carefully designed and optimized INR model architecture could further enhance performance. Finally, our current approach to model reference compression and reference initialization is relatively simplistic. Future work could focus on refining these two modules, like multi-GOP and multi-direction reference, to further improve efficiency and rate-distortion performance.